# Many-hypercube codes: High-rate quantum error-correcting codes for high-performance fault-tolerant quantum computation

### v1

@@ V Logical gate operations for C642
Partial transversal logical CNOT gates between two registers (inter-block logical CNOT gates) can be performed using two ancilla registers in the same partial |+ket|$\rangle$| + ⟩ state and full transversal logical CNOT and Hadamard gates (see Fig. 16a in Appendix M). In this case, each logical CNOT gate is performed on the corresponding two logical qubits with the same label in the two registers. To perform a logical CNOT gate on two logical qubits with different labels, we can use logical SWAP gates (see Appendix L), which allow us to perform a logical CNOT gate on an arbitrary logical-qubit pair between the two registers. Finally, logical CNOT gates in a register (intra-block logical CNOT gates) can be performed with an additional ancilla register in the logical all-zero state, logical SWAP gates, and full and partial transversal logical CNOT gates (see Fig. 16b in Appendix M).
Thus, we can implement a universal gate set (Pauli, Hadamard, phase, non-Clifford, and CNOT gates) for the C642 fault-tolerantly and perform most of them in parallel.
@@ VI Summary and outlook
We have proposed concatenated high-rate quantum error-detecting codes as a new family of high-rate quantum codes for fault-tolerant quantum computation. Since the simple code structure allows for a geometrical interpretation using hypercubes corresponding to logical qubits, we call them many-hypercube codes. The encoding rates are remarkably high, 30% and 20% for the code distances of 8 and 16, respectively. We have developed a high-performance decoder and fault-tolerant zero-state encoders dedicated to the codes. Using them, we have achieved high error thresholds: 5.6% for bit-flip errors and 0.9% for a logical CNOT gate in a circuit-level noise model. Further improvements of the decoder and encoders for the many-hypercube codes are challenging but desirable. We have also explained how to implement logical gates necessary for universal quantum computation. More efficient logical-gate implementations for the many-hypercube codes may be possible but are left for future work. Minimizing the number of logical gates and computational depth for given quantum circuits is also an important compilation problem for the many-hypercube codes.
@@ Appendix A Elementary gates
Paul $\mbox{Paul gates}:X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$ (11)
$\qquad\qquad\qquad Y=iXZ=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$ (14)
Hadamard $\mbox{Hadamard gate}:H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (17)
Phase $\mbox{Phase gate}:S=\begin{pmatrix}1&0\\ 0&i\end{pmatrix}$ (20)
@@
CNOT $\mbox{CNOT gate}:U_{\mathrm{CNOT}}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$ (28)
Appendix B The ${[\![6,4,2]\!]}$ code
The ${[\![6,4,2]\!]}$ code is one of the simplest stabilizer codes [3] and has only two stabilizer generators (check operators): $S_{Z}=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}$ and $S_{X}=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}$, which can detect an $X$ (bit-flip) error and a $Z$ (phase-flip) error, respectively. Its four logical qubits are defined by the following logical $Z$ and $X$ operators:
$Z_{L1}$ $=Z_{1}Z_{2},~{}X_{L1}=X_{2}X_{3},$ (29)
$Z_{L2}$ $=Z_{2}Z_{3},~{}X_{L2}=X_{1}X_{2},$ (30)
$Z_{L3}$ $=Z_{4}Z_{5},~{}X_{L3}=X_{5}X_{6},$ (31)
$Z_{L4}$ $=Z_{5}Z_{6},~{}X_{L4}=X_{4}X_{5}.$ (32)
Other definitions are possible [2], but we use this because of the geometrical interpretation of the code structure using hypercubes (Fig. 1). From the definition, we can easily find that transversal logical CNOT and Hadamard gates can be performed by transversal physical CNOT and Hadamard (and SWAP) gates, respectively, as shown in Figs. 5a and 5b. The logical SWAP gates between logical qubits 1 and 2, 3 and 4, and (1,2) and (3,4) can also be performed by physical SWAP gates, as shown in Fig. 5c. The logical all-zero state of the ${[\![6,4,2]\!]}$ code is the six-qubit GHZ state, i.e.,
$|0000\rangle_{L}=\frac{|000000\rangle+|111111\rangle}{\sqrt{2}}.$ (33)
Thus the zero-state encoder of the ${[\![6,4,2]\!]}$ code is given by, e.g., the quantum circuit shown in Fig. 5d. This can be generalized to an arbitrary-state encoder by adding two CNOT gates, as shown in Fig. 5e.
Figure 5: Logical gates and encoders of the ${[\![6,4,2]\!]}$ code. (a) Transversal logical CNOT gates. (b) Transversal logical Hadamard gates. (c) Logical SWAP gates. In (a–c), the left and right sides correspond to logical and physical ones, respectively. (d) Zero-state encoder. (e) Arbitrary-state encoder. In (e), additional two CNOT gates are highlighted in red.
Appendix C Error-correcting teleportation
Our quantum error correction is based on error-correcting teleportation (ECT) [44, 60, 61]. This is quantum teleportation with logical qubits, where the classical feed-forward information is determined by decoding physical-qubit measurement outcomes, as shown in Fig. 6. Assuming that the two ancilla registers are reliable sufficiently, the errors come mainly from the failure of the decoding.
Figure 6: Error-correcting teleportation. The input logical-qubit state encoded with the level-$L$ C642, $|\psi\rangle_{L}$, on the left side is teleported with two logical all-zero states to the right side. The gray bold allows show the flow of the classical information. The “D” boxes are the decoders which decode the physical-qubit measurement outcomes and provide logical-qubit ones. If a logical-qubit outcome is 1, the logical $Z$ or $X$ operator is performed on the corresponding logical qubit. The logical CNOT and Hadamard gates are performed by transversal physical CNOT and Hadamard gates, as shown in Figs. 5a and 5b.
Appendix D Hard-decision decoding of C642
@@
The hard-decision decoding of, e.g., the level-3 C642 is done as follows [44]. The measurement outcomes of physical qubits in each level-1 block and the corresponding level-1 outcomes are denoted by $\{x_{i^{\prime},j^{\prime},k^{\prime}}|k^{\prime}=1,\ldots,6\}$ and $\{x_{i^{\prime},j^{\prime},k}^{(1)}|k=1,\ldots,4\}$, respectively. All the four $x_{i^{\prime},j^{\prime},k}^{(1)}$ are set to $F$ (flag indicating an error) if the parity of the six bits, $x_{i^{\prime},j^{\prime},k^{\prime}}$, is odd according to the $Z$-stabilizer. Otherwise, they are set according to the definition of the logical $Z$ operators as follows: $x_{i^{\prime},j^{\prime},1}^{(1)}=x_{i^{\prime},j^{\prime},1}+x_{i^{\prime},j^{\prime},2}$, $x_{i^{\prime},j^{\prime},2}^{(1)}=x_{i^{\prime},j^{\prime},2}+x_{i^{\prime},j^{\prime},3}$, $x_{i^{\prime},j^{\prime},3}^{(1)}=x_{i^{\prime},j^{\prime},4}+x_{i^{\prime},j^{\prime},5}$, and $x_{i^{\prime},j^{\prime},4}^{(1)}=x_{i^{\prime},j^{\prime},5}+x_{i^{\prime},j^{\prime},6}$ (mod 2). Next, using the level-1 outcomes in each level-2 block, $\{x_{i^{\prime},j^{\prime},k}^{(1)}|j^{\prime}=1,\ldots,6\}$, we obtain the corresponding level-2 outcomes, $\{x_{i^{\prime},j,k}^{(2)}|j=1,\ldots,4\}$, as follows. If there is a single $F$ in the six $x_{i^{\prime},j^{\prime},k}^{(1)}$, e.g., $x_{i^{\prime},1,k}^{(1)}$, we can correct this as $x_{i^{\prime},1,k}^{(1)}=x_{i^{\prime},2,k}^{(1)}+x_{i^{\prime},3,k}^{(1)}+x_{i^{\prime},4,k}^{(1)}+x_{i^{\prime},5,k}^{(1)}+x_{i^{\prime},6,k}^{(1)}$ (mod 2) according to the $Z$-stabilizer. (Note that error-detecting codes can correct a located error.) Then, the four $x_{i^{\prime},j,k}^{(2)}$ are set according to the definition of the logical $Z$ operator, as above. If there is no $F$ and the parity of the six bits is even, we set the four $x_{i^{\prime},j,k}^{(2)}$ similarly. Otherwise, we set all the four $x_{i^{\prime},j,k}^{(2)}$ to $F$. Applying this decoding recursively, we finally obtain the logical-level outcomes. If it is $F$, then we randomly choose 0 or 1. For error detection used in the proposed encoders, the decoder returns $F$ (indicating detected errors) unless no $F$ is obtained throughout the decoding.
Appendix E Symbol-MAP soft-decision decoding of C642
@@
The original proposal of the soft-decision decoding [60] is not directly applicable to the C642, because this decoding is based on block-MAP (maximum a posteriori probability) decoding, but the C642 has an exponentially large number of logical-qubit values in a code block and therefore we cannot evaluate all the corresponding a posteriori probabilities efficiently. So, we modify this to a symbol-MAP version, where we calculate marginal probabilities for each logical qubit, which can be done efficiently. The symbol-MAP version of the soft-decision decoding of, e.g., the level-3 C642 is as follows. Using the prior physical-qubit error probability $p_{e}$ and the physical-qubit measurement outcomes $\{x_{i^{\prime},j^{\prime},k^{\prime}}|k^{\prime}=1,\ldots,6\}$, the a posteriori probability for a physical-qubit value $x$ is expressed as
$p_{i^{\prime},j^{\prime},k^{\prime}}(x)=p_{e}^{I(x\neq x_{i^{\prime},j^{\prime},k^{\prime}})}(1-p_{e})^{I(x=x_{i^{\prime},j^{\prime},k^{\prime}})},$ (34)
@@
where $I(A)$ is the indicator function that returns 1 if $A$ is true and otherwise 0. Then, the marginal a posteriori probability for a level-1 qubit, e.g., $\mathrm{Q}_{i^{\prime},j^{\prime},1}^{(1)}$, is given by
$p_{i^{\prime},j^{\prime},1}^{(1)}(x)=\frac{\displaystyle\sum_{x_{2}=0,1}\sum_{x_{3}=0,1}\sum_{x_{4}=0,1}R_{i^{\prime},j^{\prime}}^{(1)}(x,x_{2},x_{3},x_{4})}{\displaystyle\sum_{x_{1}=0,1}\sum_{x_{2}=0,1}\sum_{x_{3}=0,1}\sum_{x_{4}=0,1}R_{i^{\prime},j^{\prime}}^{(1)}(x_{1},x_{2},x_{3},x_{4})},$ (35)
$R_{i^{\prime},j^{\prime}}^{(1)}(x_{1},x_{2},x_{3},x_{4})$
$=\sum_{x^{\prime}_{1}=0,1}\sum_{x^{\prime}_{2}=0,1}\sum_{x^{\prime}_{3}=0,1}\sum_{x^{\prime}_{4}=0,1}\sum_{x^{\prime}_{5}=0,1}\sum_{x^{\prime}_{6}=0,1)}\prod_{k^{\prime}=1}^{6}p_{i^{\prime},j^{\prime},k^{\prime}}(x^{\prime}_{k^{\prime}})$
×$\qquad\times I\left(\sum_{k=1}^{6}x^{\prime}_{k}=0\right)I(x_{1}=x^{\prime}_{1}+x^{\prime}_{2})I(x_{2}=x^{\prime}_{2}+x^{\prime}_{3})$
@@
where the summation in the indicator functions is modulo 2. Repeating such calculations recursively, we finally obtain the marginal a posteriori probabilities for the logical-qubit values $\{x_{i,j,k}^{(3)}\}$. If $p_{i,j,k}^{(3)}(0)>0.5$, we set $x_{i,j,k}^{(3)}$ to 0 and otherwise to 1.
@@
Appendix F Level-by-level minimum distance decoding of C642
@@
The minimum distance decoding is known as one of the highest-performance decoding methods, where we search for a codeword closest to the measurement outcomes in the sense of the Hamming distance. However, there are an exponentially large number of codewords in general, hence it is usually intractable to find such a minimum-distance codeword. To solve this issue, we keep only minimum-distance codewords and discard larger-distance ones at each level from level 1 to the logical level. Since each level-1 codeword [a bit string corresponding to a level-1 computational-basis state, e.g., 000000 or 111111 for $|0000\rangle_{L}$, see Eq. (33)] consists of only six bits, we can easily select minimum-distance codewords and corresponding encoded bit strings (e.g., $0$ for $|0000\rangle_{L}$) in each level-1 block. While there is a single minimum-distance codeword (encoded bit string) when the parity of the six bits is even, there are six minimum-distance codewords (encoded bit strings) when the parity is odd. To construct level-2 codewords using the level-1 minimum-distance encoded bit strings, we first choose five level-1 blocks from the six ones of each level-2 block, and then pick one of the minimum-distance encoded bit strings from each of the five level-1 blocks. The encoded bit string of the other level-1 block is automatically determined by the parity-check condition corresponding to the $Z$-stabilizer. In general, the determined encoded bit string is not included in the minimum-distance ones of this block. So we evaluate the distance of this encoded bit string, and then obtain the distance of the level-2 block by summing the distances of the six level-1 blocks. We select the minimum-distance codewords (encoded bit strings) of the level-2 block among all the choices of the five level-1 blocks and their encoded bit strings, and keep only them as minimum-distance candidates. This is our strategy to utilize the minimum-distance candidates while satisfying the parity-check condition. Doing this selection recursively, we finally obtain the minimum-distance candidates at the logical level. If we have multiple candidates at the logical level, we randomly choose one of them. For error detection used in the proposed encoders, the decoder returns $F$ unless all the numbers of the candidates at the levels from $L_{D}$ to $L$ are one, where $L$ is the logical level and $L_{D}$ ($\leq L$) is chosen appropriately.
In the following, we explain our proposed decoding method in more detail using flowcharts. As an example, we mainly focus on the level-3 case. Figure 7 shows the whole process of the decoding in the level-3 case. The subroutines Sub1 and Sub2 are explained above. So in the following, we explain the subroutine Sub3 in detail.
Figure 8 shows the process of Sub3. In the subroutine Sub31, we first construct level-3 codewords by determining the encoded bit string of the $b_{2}$th level-2 block ($b_{2}=1,\ldots,6$) using the minimum-distance encoded bit strings of the other five level-2 blocks determined in Sub2 according to the parity-check condition ($Z$-stabilizer condition). In general, the level-2 encoded bit string determined here is not included in the minimum-distance encoded bit strings of this block determined in Sub2. So in the subroutine Sub32 explained later, we evaluate the distance of this encoded bit string. Then we evaluate the distances of the level-3 codewords by summing the distances of their six level-2 blocks. We finally select level-3 encoded bit strings corresponding to the level-3 codewords with minimum distance.
Figure 9 shows the details of Sub31 when $b_{2}=6$, where $N_{1}$ to $N_{5}$ denote the numbers of the minimum-distance candidates of the first to fifth level-2 blocks, respectively. If the total number of candidates, $N_{1}N_{2}N_{3}N_{4}N_{5}$, is larger than a preset threshold $N_{\mathrm{th}3}$, we reduce the number by randomly choosing one of the candidates from the level-2 block with the largest $N$, in order to keep the decoding time short. In this work, we set $N_{\mathrm{th}3}$ to $10^{5}$. Then we determine the encoded bit string of the sixth level-2 block using the $n_{1}$th to $n_{5}$th minimum-distance encoded bit strings of the other five level-2 blocks, respectively, according to the parity-check condition. More concretely, the encoded bit string of the sixth level-2 block is determined as follows:
$x_{6,j,l}^{(2)}=x_{1,j,l}^{(2)}+x_{2,j,l}^{(2)}+x_{3,j,l}^{(2)}+x_{4,j,l}^{(2)}+x_{5,j,l}^{(2)},$ (37)
where the summation is modulo 2 and $j=1,\ldots,4,l=1,\ldots,4$. The distance of the encoded bit string determined here is evaluated in the subroutine Sub32. We finally construct the level-3 codeword using the encoded bit strings of the six level-2 blocks and evaluate the level-3 distance by summing the level-2 distances.
Figure 10 shows the subroutine Sub32 when $b_{2}=6$, where we evaluate the distance of the sixth level-2 block determined in Sub31, as mentioned above. In Fig. 10, $M_{1}$ to $M_{6}$ denote the numbers of the minimum-distance candidates of the first to sixth level-1 blocks, respectively, in the sixth level-2 block. If $M_{1}+M_{2}+M_{3}+M_{4}+M_{5}+M_{6}$ is larger than a preset threshold $M_{\mathrm{th}2}$, we reduce the number by randomly choosing one of the candidates from the level-1 block with the largest $M$. In this work, we set $M_{\mathrm{th}2}$ to 6. Then in the subroutine Sub33, we evaluate the distance of the sixth level-2 block when the minimum-distance candidates of the $b_{1}$th level-1 block ($b_{1}=1,\ldots,6$) are used. We finally select the minimum distance among all the choices of the level-1 blocks and their minimum-distance candidates.
Figure 11 shows the subroutine Sub33 when $b_{1}=1$. We first determine the encoded bit strings of the five level-1 blocks other than the first level-1 block in the sixth level-2 block using the m1th minimum-distance encoding bit string of the first level-1 block determined in Sub1 and the level-2 encoding bit string determined in Sub31 according to the parity-check condition and the definition of the encoded $Z$ operator. More concretely, the encoded bit strings of the five level-1 blocks are determined as follows:
$x_{6,2,l}^{(1)}$ $=x_{6,1,l}^{(1)}+x_{6,1,l}^{(2)},$ (38)
$x_{6,3,l}^{(1)}$ $=x_{6,2,l}^{(1)}+x_{6,2,l}^{(2)},$ (39)
$x_{6,5,l}^{(1)}$ $=x_{6,2,l}^{(1)}+x_{6,1,l}^{(2)}+x_{6,2,l}^{(2)}+x_{6,3,l}^{(2)}+x_{6,4,l}^{(2)},$ (40)
$x_{6,4,l}^{(1)}$ $=x_{6,5,l}^{(1)}+x_{6,3,l}^{(2)},$ (41)
$x_{6,6,l}^{(1)}$ $=x_{6,5,l}^{(1)}+x_{6,4,l}^{(2)},$ (42)
where the summation is modulo 2 and $l=1,\ldots,4$. We obtain the distance of the sixth level-2 block by summing the distances of the six level-1 blocks.
The decoding of the level-4 C642 is done in a similar manner to the above level-3 case. In the level-4 case, we set the level-4 threshold $N_{\mathrm{th}4}$ corresponding to the above level-3 threshold $N_{\mathrm{th}3}$ to the same value $10^{5}$ and the level-3 threshold $M_{\mathrm{th}3}$ corresponding to the above level-2 threshold $M_{\mathrm{th}2}$ to 12.
Appendix G Simulation of the bit-flip error model

### v3

@@ V Logical gate operations
Partial transversal logical CNOT gates between two registers (inter-block logical CNOT gates) can be performed using two ancilla registers in the same partial |+ket|$\rangle$| + ⟩ state and full transversal logical CNOT and Hadamard gates, as shown in Fig. 9A. In this case, each logical CNOT gate is performed on the corresponding two logical qubits with the same label in the two registers. Using logical SWAP gates explained above, we can perform at least a logical CNOT gate on an arbitrary logical-qubit pair between the two registers. Finally, logical CNOT gates in a register (intra-block logical CNOT gates) can be performed with an ancilla register in the logical all-zero state, logical SWAP gates, and inter-block full and partial transversal logical CNOT gates, as shown in Fig. 9B.
Thus, we can implement a universal gate set for the many-hypercube codes fault-tolerantly and perform most of them in parallel. For qLDPC codes, we can, in principle, achieve parallel execution of logical gates in a similar manner, but then logical ancilla states required for it are difficult to prepare, which may need another high-threshold code such as concatenated Steane codes [18]. On the other hand, for high-rate concatenated codes, such as the many-hypercube codes and the concatenated quantum Hamming codes [44], the required ancilla states themselves are encoded with the same high-rate concatenated codes and therefore relatively easy to prepare. This is the advantage of high-rate concatenated codes over qLDPC codes. To overcome the ancilla-preparation issue, a current standard approach to logical gates for qLDPC codes uses the teleportation between a logical qubit in a qLDPC quantum memory and another logical qubit encoded with, e.g., the surface code [39, 40], where logical gates are performed on the surface-code logical qubits. Then, however, parallel execution of $N$ logical gates requires at least $N$ surface-code logical qubits, leading to large overheads for large $N$. Thus parallel execution of logical gates for qLDPC codes is still challenging.
@@ VI Summary and outlook
We have proposed concatenated high-rate quantum error-detecting codes as a new family of high-rate quantum codes for fault-tolerant quantum computation. Since the simple code structure allows for a geometrical interpretation using hypercubes corresponding to logical qubits, we call them many-hypercube codes. The encoding rates are remarkably high, 30% and 20% (64 and 256 logical qubits are encoded into 216 and 1296 physical qubits, respectively) for the code distances of 8 and 16, respectively. We have developed a high-performance decoder and fault-tolerant zero-state encoders dedicated to the codes. Using them, we have achieved high error thresholds: 5.6% for bit-flip errors and 0.9% for a logical CNOT gate in a circuit-level noise model. Further improvements of the decoder and encoders for the many-hypercube codes are challenging but desirable. We have also explained how to implement logical gates necessary for universal quantum computation. More efficient logical-gate implementations for the many-hypercube codes may be possible but are left for future work. Minimizing the number of logical gates and computational depth for given quantum circuits is also an important compilation problem for the many-hypercube codes.
@@ Appendix A Elementary gates
Identity $\mbox{Identity gate}:I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},$
Paul $\mbox{Paul gates}:X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$
$\qquad\qquad\qquad Y=iXZ=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$
Hadamard $\mbox{Hadamard gate}:H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$
Phase $\mbox{Phase gate}:S=\begin{pmatrix}1&0\\ 0&i\end{pmatrix}$
@@
CNOT $\mbox{CNOT gate}:U_{\mathrm{CNOT}}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$
Appendix B Hard-decision decoding of the many-hypercube codes
@@
The hard-decision decoding of, e.g., the level-3 many-hypercube code is done as follows [65]. The measurement outcomes of physical qubits in each level-1 block and the corresponding level-1 outcomes are denoted by $\{x_{i,j,k}|i=1,\ldots,6\}$ and $\{x_{i^{\prime},j,k}^{(1)}|i^{\prime}=1,\ldots,4\}$, respectively. All the four $x_{i^{\prime},j,k}^{(1)}$ are set to $F$ (flag indicating an error) if the parity of the six bits, $x_{i,j,k}$, is odd according to the $Z$-stabilizer in Eqs. (1)–(4). Otherwise, they are set according to the definition of the logical $Z$ operators as follows: $x_{1,j,k}^{(1)}=x_{1,j,k}+x_{2,j,k}$, $x_{2,j,k}^{(1)}=x_{2,j,k}+x_{3,j,k}$, $x_{3,j,k}^{(1)}=x_{4,j,k}+x_{5,j,k}$, and $x_{4,j,k}^{(1)}=x_{5,j,k}+x_{6,j,k}$ (mod 2). Next, using the level-1 outcomes in each level-2 block, $\{x_{i^{\prime},j,k}^{(1)}|j=1,\ldots,6\}$, we obtain the corresponding level-2 outcomes, $\{x_{i^{\prime},j^{\prime},k}^{(2)}|j^{\prime}=1,\ldots,4\}$, as follows. If there is a single $F$ in the six $x_{i^{\prime},j,k}^{(1)}$, e.g., $x_{i^{\prime},1,k}^{(1)}$, we can correct this as $x_{i^{\prime},1,k}^{(1)}=x_{i^{\prime},2,k}^{(1)}+x_{i^{\prime},3,k}^{(1)}+x_{i^{\prime},4,k}^{(1)}+x_{i^{\prime},5,k}^{(1)}+x_{i^{\prime},6,k}^{(1)}$ (mod 2) according to the $Z$-stabilizer. (Note that error-detecting codes can correct a located error.) Then, the four $x_{i^{\prime},j^{\prime},k}^{(2)}$ are set according to the definition of the logical $Z$ operator, as above. If there is no $F$ and the parity of the six bits is even, we set the four $x_{i^{\prime},j^{\prime},k}^{(2)}$ similarly. Otherwise, we set all the four $x_{i^{\prime},j^{\prime},k}^{(2)}$ to $F$. Applying this decoding recursively, we finally obtain the logical-level outcomes. If it is $F$, then we randomly choose 0 or 1. For error detection used in the proposed encoders, the decoder returns $F$ (indicating detected errors) unless no $F$ is obtained throughout the decoding.
Appendix C Symbol-MAP decoding of the many-hypercube codes
@@
In the symbol-MAP decoding, we calculate marginal probabilities for each logical qubit, which can be done efficiently. For example, the symbol-MAP decoding of the level-3 many-hypercube code is as follows. Using the prior physical-qubit error probability $p_{e}$ and the physical-qubit measurement outcomes $\{x_{i,j,k}|i=1,\ldots,6\}$, the a posteriori probability for a physical-qubit value $x$ is expressed as
$p_{i,j,k}(x)=p_{e}^{I(x\neq x_{i,j,k})}(1-p_{e})^{I(x=x_{i,j,k})},$ (20)
@@
where $I(A)$ is the indicator function that returns 1 if $A$ is true and otherwise 0. Then, the marginal a posteriori probability for a level-1 qubit, e.g., $\mathrm{Q}_{1,j,k}^{(1)}$, is given by
$p_{1,j,k}^{(1)}(x)=\frac{\displaystyle\sum_{x_{2}=0,1}\sum_{x_{3}=0,1}\sum_{x_{4}=0,1}R_{j,k}^{(1)}(x,x_{2},x_{3},x_{4})}{\displaystyle\sum_{x_{1}=0,1}\sum_{x_{2}=0,1}\sum_{x_{3}=0,1}\sum_{x_{4}=0,1}R_{j,k}^{(1)}(x_{1},x_{2},x_{3},x_{4})},$ (21)
$R_{j,k}^{(1)}(x_{1},x_{2},x_{3},x_{4})$
$=\sum_{x^{\prime}_{1}=0,1}\sum_{x^{\prime}_{2}=0,1}\sum_{x^{\prime}_{3}=0,1}\sum_{x^{\prime}_{4}=0,1}\sum_{x^{\prime}_{5}=0,1}\sum_{x^{\prime}_{6}=0,1}\prod_{i=1}^{6}p_{i,j,k}(x^{\prime}_{i})$
×$\qquad\times I\left(\sum_{i=1}^{6}x^{\prime}_{i}=0\right)I(x_{1}=x^{\prime}_{1}+x^{\prime}_{2})I(x_{2}=x^{\prime}_{2}+x^{\prime}_{3})$
@@
where the summation in the indicator functions is modulo 2. Repeating such calculations recursively, we finally obtain the marginal a posteriori probabilities for the logical-qubit values $\{x_{i^{\prime},j^{\prime},k^{\prime}}^{(3)}\}$. If $p_{i^{\prime},j^{\prime},k^{\prime}}^{(3)}(0)>0.5$, we set $x_{i^{\prime},j^{\prime},k^{\prime}}^{(3)}$ to 0 and otherwise to 1.
@@
Appendix D Level-by-level minimum distance decoding of the many-hypercube codes
@@
The minimum distance decoding is known as one of the highest-performance decoding methods, where we search for a codeword closest to the measurement outcomes in the sense of the Hamming distance. However, there are an exponentially large number of codewords in general, hence it is usually intractable to find such a minimum-distance codeword. To solve this issue, we keep only minimum-distance codewords and discard larger-distance ones at each level from level 1 to the logical level. Since each level-1 codeword [a bit string corresponding to a level-1 computational-basis state, e.g., 000000 or 111111 for $|0000\rangle_{L}$, see Eq. (5)] consists of only six bits, we can easily select minimum-distance codewords and corresponding encoded bit strings (e.g., $0$ for $|0000\rangle_{L}$) in each level-1 block. While there is a single minimum-distance codeword (encoded bit string) when the parity of the six bits is even, there are six minimum-distance codewords (encoded bit strings) when the parity is odd. To construct level-2 codewords using the level-1 minimum-distance encoded bit strings, we first choose five level-1 blocks from the six ones of each level-2 block, and then pick one of the minimum-distance encoded bit strings from each of the five level-1 blocks. The encoded bit string of the other level-1 block is automatically determined by the parity-check condition corresponding to the $Z$-stabilizer. In general, the determined encoded bit string is not included in the minimum-distance ones of this block. So we evaluate the distance of this encoded bit string, and then obtain the distance of the level-2 block by summing the distances of the six level-1 blocks. We select the minimum-distance codewords (encoded bit strings) of the level-2 block among all the choices of the five level-1 blocks and their encoded bit strings, and keep only them as minimum-distance candidates. This is our strategy to utilize the minimum-distance candidates while satisfying the parity-check condition. Doing this selection recursively, we finally obtain the minimum-distance candidates at the logical level. If we have multiple candidates at the logical level, we randomly choose one of them. See Appendix I for more details of this decoding method.
For error detection used in the proposed encoders, the decoder returns $F$ unless all the numbers of the candidates at the levels from $L_{D}$ to $L$ are one, where $L$ is the logical level and $L_{D}$ ($\leq L$) is chosen appropriately. The most stringent condition for error detection is given by $L_{D}=1$. To increase the acceptance probabilities at error-detection gadgets in the level-4 zero-state encoder, we relax the condition by setting $L_{D}$ as $L_{D}=2$.
Appendix E Simulation of the bit-flip error model